Add:

29+(-42)+510+(-397)+52(-792)
Help me please. I do not understand am i supposed to subtract all the ones in parenthisis then add the rest, then subtract the total parenthisis from the total of the ones not in parenthesis?

Answer by checkley71(8403)

SEQUENCE DOEN'T MATTER IN THIS EXAMPLE.
WHAT I WOULD DO IS ADD ALL THE + NUMBERS TOGETHER THUS:
29+510+52=591
THEN ADD UP THE - NUMBERS:
-42-397-792=-1,231
NOW FIND THE DIFFERENCE.
591-1,231=-640 ANSWER.
YOU COULD ALSO COMBINE THEN AS THEY ARE ARRANGED:
(29-42)=(-13+510)=(497-397)=(100+52)=(152-792)=-640